OpenStudy (anonymous):

Variable y varies directly with variable x, and y = 18 when x = 4. What is the constant of variation in decimal form?

OpenStudy (welshfella):

direct variation is y = kx where k is a constant to be found

OpenStudy (welshfella):

Plug in your 2 values and find the value of k

OpenStudy (anonymous):

Huh i'm not really good at math can you dumb it down a little

OpenStudy (anonymous):

2*9=18

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

so whats the decimal form

OpenStudy (anonymous):

0.9?

OpenStudy (welshfella):

sorry - I lost it then y = kx y = 18 when x = 4 so plugging in theses values 18 = k*4 now to get k on its own you divide each side of the eqaution by 4:- 18/4 = 4k / 4 18/4 = k

OpenStudy (anonymous):

so whats k? I'm confuzed

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

k is the constant of variation which you are trying to find all you need to do now is divide 18 by 4 and you have it k = 18/4 = ?

OpenStudy (anonymous):

4.5?

OpenStudy (welshfella):

yes